OpenStudy (marziajackson$):

What is the value of the expression 2/4^-3 ?

OpenStudy (michele_laino):

hint: \[\Large {4^{ - 3}} = \frac{1}{{{4^3}}} = \frac{1}{{4 \times 4 \times 4}} = \frac{1}{{64}}\]